OpenStudy (anonymous):

Algebra II Help! Please! I'm not asking for an answer, I just need help! Let f(x) = 4x^2 + x + 1 and g(x) = x^2 – 2. Find g(f(x)). Show each step of your work.

OpenStudy (misty1212):

HI!!

Vocaloid (vocaloid):

g(f(x)) means look at g(x) = x^2 - 2, and replace the "x" with f(x), which equals 4x^2 + x + 1

OpenStudy (misty1212):

work from the inside out \[g(f(x))=g(4x^2+x+1)\] then since \[g(\spadesuit)=\spadesuit^2-2\] you have \[g(f(x))=g(4x^2+x+1)=(4x^2+x+1)^2-2\]

OpenStudy (anonymous):

I'm sorry, I'm not understanding? I'm normally decent with math but I have a difficult time with functions. :(

Vocaloid (vocaloid):

well, we know that g(x) = x^2 - 2, right? if we replace x with f(x) on both sides, what do we get?

Vocaloid (vocaloid):



Vocaloid (vocaloid):

does it make a little more sense now?

OpenStudy (anonymous):

Kind of...

OpenStudy (anonymous):

I think I got it. Thank you both for your help.